Among 250 employees, of the local office of an international insurance company, 182 are

whites, 51 are blacks and the rest are Orientals. If 5% margin of error is used,

5. How many employees will be chosen for a case study as samples?

6. How many Orientals will be chosen?

7. How many whites will be chosen?


1
Expert's answer
2022-01-25T14:30:06-0500

Margin of error = 5%

e= 0.05

Confidence level = 100-5= 95%

Population size, N = 250

Since here we don't know the population proportion so, p= 0.5 is assumed.

Now we find Z at 95% confidence level using the normal distribution table and that is

Z= 1.96 at 95% confidence level

5. The formula for sample size:

"n= \frac{\frac{Z^2p(1-p)}{e^2}}{1+\frac{Z^2p(1-p)}{Ne^2}} \\

n= \frac{\frac{1.96^2 \times 0.5(1-0.5)}{0.5^2}}{1+\frac{1.96^2 \times 0.5(1-0.5)}{2050 \times 0.05^2}} \\

n=151.44 = 152"

6. In the population , the number of orientals = 250-182-51=17

The proportion of orientals

"p= \frac{17}{250} = 0.068"

Substituting this value of p in the above formula we get the number of orientals in the sample size of 152.

"n_{Oriental} = \frac{\frac{1.96^2 \times 0.068(1-0.068)}{0.05^2}}{1 + \frac{1.96^2 \times 0.068(1-0.068)}{250 \times 0.05^2}} \\

= 70.08 = 71"

7. Number of whites in the population = 182

Proportion of whites

"p= \frac{182}{250}= 0.728"

On substituting this value of p in the above formula we get the number of whites in the sample.

"n_{Whites} = \frac{\frac{1.96^2 \times 0.728(1-0.728)}{0.05^2}}{1 + \frac{1.96^2 \times 0.728(1-0.728)}{250 \times 0.05^2}} \\

= 137.24 = 138"